if the round object is a ring, using I=mR^2, what is the fraction of rotational kinetic energy over the total kinetic energy?
1 : 1/4
2 : 1/3
3 : 1/2
4 : 2/3
5 : 1
1
Expert's answer
2020-06-09T13:18:53-0400

The total kinetic energy of a disk moving with a speed "v" is the sum of kinetic energy of linear (L) and rotational (R) motion:


"KE=KE_L+KE_R,\\
KE_L=\frac{mv^2}{2},\space KE_R=\frac{I\omega^2}{2}=\frac{[mR^2][v/R]^2}{2}=\frac{mv^2}{2}.\\
KE=\frac{mv^2}{2}+\frac{mv^2}{2}=mv^2."

Therefore, the fraction of rotational kinetic energy over the total kinetic energy is 1/2.